मराठी

Two Dice Are Thrown Together. the Probability that Neither They Show Equal Digits Nor the Sum of Their Digits is 9 Will Be (A) 13/15 (B) 13/18 (C) 1/9 (D) 8/9

Advertisements
Advertisements

प्रश्न

Two dice are thrown together. The probability that neither they show equal digits nor the sum of their digits is 9 will be

पर्याय

  •  13/15

  •  13/18

  •  1/9

  •  8/9

     
MCQ
Advertisements

उत्तर

13/18

When two dices are thrown, there are (6 × 6) = 36 outcomes.
The set of all these outcomes is the sample space is given by
S = (1, 1) , (1, 2), (1, 3), (1, 4), (1, 5), (1, 6)
      (2, 1) , (2, 2), (2, 3), (2, 4), (2, 5), (2, 6)
      (3, 1) , (3, 2), (3, 3), (3, 4), (3, 5), (3, 6)
      (4, 1) , (4, 2), (4, 3), (4, 4), (4, 5), (4, 6)
      (5, 1) , (5, 2), (5, 3), (5, 4), (5, 5), (5, 6)
      (6, 1) , (6, 2), (6, 3), (6, 4), (6, 5), (6, 6)
∴ n(S) = 36
Let E be the event of getting the digits which are neither equal nor give a total of 9.
Then E' = event of getting either a doublet or a total of 9
Thus, E' = {{1, 1), (2, 2), (3, 3), (4, 4), (5, 5), (6, 6), (3, 6), (4, 5), (5, 4), (6, 3)}
i.e. n(E') = 10
P(E') = \[\frac{n\left( E' \right)}{n\left( S \right)} = \frac{10}{36} = \frac{5}{18}\]

Hence, required probability P(E) = 1- P(E')
                                                   = \[1 - \frac{5}{18} = \frac{13}{18}\]

 
shaalaa.com
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 33: Probability - Exercise 33.6 [पृष्ठ ७१]

APPEARS IN

आर.डी. शर्मा Mathematics [English] Class 11
पाठ 33 Probability
Exercise 33.6 | Q 8 | पृष्ठ ७१

व्हिडिओ ट्यूटोरियलVIEW ALL [1]

संबंधित प्रश्‍न

Which of the following can not be valid assignment of probabilities for outcomes of sample space S = {ω1, ω2,ω3,ω4,ω5,ω6,ω7}

Assignment ω1 ω2 ω3 ω4 ω5 ω6 ω7
(a) 0.1 0.01 0.05 0.03 0.01 0.2 0.6
(b) `1/7` `1/7` `1/7` `1/7` `1/7` `1/7` `1/7`
(c) 0.1 0.2 0.3 0.4 0.5 0.6 0.7
(d) –0.1 0.2 0.3 0.4 -0.2 0.1 0.3
(e) `1/14` `2/14` `3/14` `4/14` `5/14` `6/14` `15/14`

A fair coin with 1 marked on one face and 6 on the other and a fair die are both tossed. Find the probability that the sum of numbers that turn up is (i) 3 (ii) 12


There are four men and six women on the city council. If one council member is selected for a committee at random, how likely is it that it is a woman?


Check whether the following probabilities P(A) and P(B) are consistently defined

P(A) = 0.5, P(B) = 0.4, P(A ∪ B) = 0.8


Fill in the blank in following table:

P(A) P(B) P(A ∩ B) P(A ∪ B)
0.5 0.35 .... 0.7

If 4-digit numbers greater than 5,000 are randomly formed from the digits 0, 1, 3, 5, and 7, what is the probability of forming a number divisible by 5 when, the digits are repeated?


Two unbiased dice are thrown. Find the probability that the total of the numbers on the dice is greater than 10.

 

Two unbiased dice are thrown. Find the probability that  neither a doublet nor a total of 8 will appear


Two unbiased dice are thrown. Find the probability that the sum of the numbers obtained on the two dice is neither a multiple of 2 nor a multiple of 3


A bag contains 8 red, 3 white and 9 blue balls. If three balls are drawn at random, determine the probability that all the three balls are blue balls 


In a lottery, a person chooses six different numbers at random from 1 to 20, and if these six numbers match with six number already fixed by the lottery committee, he wins the prize. What is the probability of winning the prize in the game?


Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(ii)
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]
\[\frac{1}{7}\]

Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(iii) 0.7 0.06 0.05 0.04 0.03 0.2 0.1

Which of the cannot be valid assignment of probability for elementary events or outcomes of sample space S =  {w1, w2, w3, w4, w5, w6, w7}:

Elementary events: w1 w2 w3 w4 w5 w6 w7
(iv)
\[\frac{1}{14}\]
\[\frac{2}{14}\]
\[\frac{3}{14}\]
\[\frac{4}{14}\]
\[\frac{5}{14}\]
\[\frac{6}{14}\]
\[\frac{15}{14}\]

In a single throw of three dice, find the probability of getting the same number on all the three dice.


A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability that: all 10 are defective


A box contains 100 bulbs, 20 of which are defective. 10 bulbs are selected for inspection. Find the probability thatat least one is defective


Three squares of chessboard are selected at random. The probability of getting 2 squares of one colour and other of a different colour is ______.


One mapping (function) is selected at random from all the mappings of the set A = {1, 2, 3, ..., n} into itself. The probability that the mapping selected is one to one is ______.


Four candidates A, B, C, D have applied for the assignment to coach a school cricket team. If A is twice as likely to be selected as B, and B and C are given about the same chance of being selected, while C is twice as likely to be selected as D, what are the probabilities that C will be selected?


Four candidates A, B, C, D have applied for the assignment to coach a school cricket team. If A is twice as likely to be selected as B, and B and C are given about the same chance of being selected, while C is twice as likely to be selected as D, what are the probabilities that A will not be selected?


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine P(A ∪ B)


The accompanying Venn diagram shows three events, A, B, and C, and also the probabilities of the various intersections (for instance, P(A ∩ B) = .07). Determine Probability of exactly one of the three occurs


A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all the three balls are white


A bag contains 8 red and 5 white balls. Three balls are drawn at random. Find the probability that all the three balls are red


If the letters of the word ASSASSINATION are arranged at random. Find the probability that four S’s come consecutively in the word


If the letters of the word ASSASSINATION are arranged at random. Find the probability that all A’s are not coming together


In a non-leap year, the probability of having 53 tuesdays or 53 wednesdays is ______.


Three numbers are chosen from 1 to 20. Find the probability that they are not consecutive ______.


While shuffling a pack of 52 playing cards, 2 are accidentally dropped. Find the probability that the missing cards to be of different colours ______.


Seven persons are to be seated in a row. The probability that two particular persons sit next to each other is ______.


6 boys and 6 girls sit in a row at random. The probability that all the girls sit together is ______.


The probability that a person visiting a zoo will see the giraffee is 0.72, the probability that he will see the bears is 0.84 and the probability that he will see both is 0.52.


The probability that a student will pass his examination is 0.73, the probability of the student getting a compartment is 0.13, and the probability that the student will either pass or get compartment is 0.96.


The probabilities that a typist will make 0, 1, 2, 3, 4, 5 or more mistakes in typing a report are, respectively, 0.12, 0.25, 0.36, 0.14, 0.08, 0.11. 


The sum of probabilities of two students getting distinction in their final examinations is 1.2


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×